You are planning to withdraw $100 in Year 1, $150 in Year 3, and $200 in Year 5, At a 5% interest rate. what is the present worth of these withdrawals?
Answer #1

Answer

Present value of amount A is given by:

PV = A/(1 + r)n

where n= time , A = Amount and r = interest rate = 5% = 0.05

Hence Present value of 100 that has been withrawn after 1 year is given by:

PV = 100/(1 + 0.05)

Present value of 150 that has been withdrawn after 3 year is given by:

PV = 150/(1 + 0.05)3

Present value of 200 that has been withdrawn after 5 years is given by:

PV = 200/(1 + 0.05)5

Hence Net present worth of all withdrawals are :

NPW = 100/(1 + 0.05) + 150/(1 + 0.05)3 + 200/(1 + 0.05)5 = 381.52

Hence Net Present worth of these withdrawals = $381.52